Home‎ > ‎Reports‎ > ‎Crew103‎ > ‎

day08


Crew 103 Log Book for April 16, 2011

Photos of the day





Commander's Check-In Report

Anne-Sophie Kalbfleisch Reporting

Crew Physical Status: Everybody's fine Time Departed/Returned from EVA:

11h30-12h

Brief Narrative of Field Mission Results:

The aim of the EVA was to install the antenna outside as a part of the

radiotelescope.

We spent almost the whole day with our experiments with the algae. We

made some measurements in order to determine the optical density of

our algae solution. Each measurement was made 5 time in order to get a

good average and this took a lot of time. After that, we assembled

the tubes and pipes but until now we have had some problems because

the junctions weren't airtight.

EVA Data/Interpretations:

Until now there were no results because there was no solar eruption.

For more details see the EVA report.

Engineering/Hab Maintenance:

Nothing to mention

Plans for Tomorrow:

We will continue our experiments with the algae's. Perhaps one team

will continue the analysis of the surroundig soils.

Inventory:/

Support Requested:

Nothing for the moment

Miscellaneous:

We are trying to cook tofu and make it tasty... difficult job :)


Commander's Report

Anne-Sophie Kalbfleisch Reporting

This morning W. de Groot and G. Van Roy made an EVA in order to install the antennae for the radio-telescope. They also connected the spectrom. The ensemble will be used for detecting solar winds (see our website for more information). In the afternoon we started our experiments with the algae. Firstly, we determined the optical density of our sample (as well as the optical signal of the solution without algae that will be substracted from the original signal). We then constructed the ensemble which consisted of two tubes, one filled with water to bring some water vapor in the air for the algae, the other filled with our solution. Both are connected and will be placed in a dark place with light coming from our LED ensemble. We encountered one major problem: our tubes weren't airtight. We will try to repair it tomorrow.


Engineering Report

Thibault Windels Reporting

Generator/Electricity:

Honey (Diesel Generator) (Used? Yes/No): yes

Wendy (Gas Generator) (Used? Yes/No): no

Battery Charge Level: 10,3 volts

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0,5

Propane Tank (Full/0.75/0.5/0.25/Empty): 0,39

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 10 galons

Notes/Comments: /


Oil:

Kitty Oil Quantity (# of Quarts in storage): the previous commander told us not to check the oil quantity of Kitty because we have to turn the all generator off!

Honey Oil Quantity (# of Quarts in storage): /

ATV Oil Quantity (# of Quarts in storage): full

Car Oil Quantity (# of Quarts in storage): /


Water Status:

Outside Potable Water Tank Level (inches from bottom): 14,5

Trailer Potable Water Tank Level (inches from bottom): 12,5

Hab Potable Water Tank Level (inches from bottom): 9

Potable Water Meter Reading: 45181,2

In to GreenHab Meter Reading: /

Notes/Comments: /


GreenHab:

Condition of plants in Tank 1: /

Condition of plants in Tank 2: /

Condition of plants in Tank 3: /

Greenhab Notes/Comments: The ventilation of the green hab is on . The thermostat is at 50°F and the temperature inside is 82°F


Observatory:

Telescope Used? (Yes/No): no

Observatory Notes/Comments: there's no telescope


Transportation:


Hab Car (NOT RENTAL)

Used (yes/no): no

Oil Checked (yes/no): no

Tire Status: ok

Odometer at end of day: 124250,5

Notes/Comments on Hab Car: /


ATV 1 (Opportunity):

Used: no

Oil Checked: no

Fuel consumed : 0

Tire Status: ok

Comments: /


ATV 2 (Spirit):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status: ok

Comments: /


ATV 3 (Viking I):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status: ok

Comments: /


ATV 4 (Viking II):

Used: no

Oil Checked: no

Fuel Consumed: 0

Tire Status:

Comments: does not work


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 60

Actual temperature upstairs (Farenheit): 79

Thermostat setting downstairs (Farenheit): /

Actual temperature downstairs (Farenheit): 76


Computers/Networking Infrastructure:

Notes/comments on internet/computers?: internet is still running well


General Engineering Notes/Comments: I put a bulb in the EVA room. I have one question about the cams, where is the 'outside south cam'? I'd like to check this cam because the pic is always very dark...


Science Report (Biology)

Wouter de Groot Reporting

Today we worked all the day on the assembly of the experiment in aqueous environment and took our first pictures of the algae on solid substrate.

As you will see in one of the picture Jonas will send you, we have installed the assembly in one of the drawers of the Lab. The assembly is composed of a stative where three LEDs are fixed, we chose the blue ones to start. Then just at the right of the stative is the tube containing the algae solution, this tube is linked to another one containing only water which is linked to an aquarium pump. The pump blows air into the water tube where it will absorb water vapor, then this air in injected into the algae tube to feed the algae in carbon dioxide.


The reason why the picture doesn't show the experiment in action is because at the moment, we didn't succeed in making the top of the fist tube completely airtight, so we can't use it so far and instead injected the air directly into the algae tube.

To make measurements, we use a 40 watt halogen lamp to light the tube and at the other side we have a 750nm optical filter in front of a silicate photo-diode and we measure the transmitted light intensity. The higher the algae concentration will be, the lower will be the intensity detected. We already started the measurements.


EVA 5 Report

Gregory Van Roy Reporting

Time: 11.30 - 12 am

Crew Members: Wouter De Groot and Gregory Van Roy Site

Location: Front of the station

Transit Mode: by foot

Objectives: Install the antenna of the radio-telescope

Results: The antenna and the spectrum analysis work well

Lessons Learned: None


Chef's Report   

Gregory Van Roy Reporting

1a) Today was

A cooking day

1b) Was today a special day celebrated at the main meal with special food or activity?

We arrived here one week ago, we will take some pepperoni and nuts for

aperitif.

2)  List the foods served at today's main meal (usually the evening meal),

giving full names of each.

BeerBreak Soup Creamy Potato,Tortillas with cheddar cheese, tomato sauce, sweetcorn and onion and Strawberry drink mix

3) List any main meal foods not finished by the crew (leftovers stored for

later use or discarded)

Sweetcorn : 1/2 of the pack left

Cheddar : 1/3 of the pack left

Onion : 1/2 of the pack left

4)  Recipes for anything more complicated than following package

directions exactly, or re-hydrating.

Chocolate cake : 2 cups sugar, 2 cups flour, 1 TSP Baking Soda, 1/2 TSP

salt, 1 cup of water, 1/2 cup oil, 8 TBSP butter, 1/4 cup cocoa, 2 eggs,

1/2 cup milk

5)  Lessons learned  (if any)

Tofu has a bad taste...

6) Comments/questions for Kim and Jean, the food study investigators

We decided to use milk only for Quaker and bread because of the low

reserve we have.